A CHEMIST HAS TWO SOLUTIONS OF H2SO4. ONE HAS 40% CONCENTRATION AND THE OTHER HAS 25% CONCENTRATION. HOW MANY LITERS OF EACH SOLUTION MUST BE MIXED TO OBTAIN 50 LITERS OF A 27% SOLUTION? ROUND TO THE NEAREST TENTH IF POSSIBLE
Answer by stanbon(75887)

A CHEMIST HAS TWO SOLUTIONS OF H2SO4. ONE HAS 40% CONCENTRATION AND THE OTHER HAS 25% CONCENTRATION. HOW MANY LITERS OF EACH SOLUTION MUST BE MIXED TO OBTAIN 50 LITERS OF A 27% SOLUTION? ROUND TO THE NEAREST TENTH IF POSSIBLE
------
Equation:
active + active = active
0.40x + 0.25(50-x) = 0.27*50
---
Multiply thru by 100 to get:
40x + 25*50 - 25x = 27*50
15x = 2*50
x = 100/15 = 6 2/3 liters (amt. of 40% solution needed)
-----
50-x = 43 1/3 liters (amt. of 25% solution needed)
=================================
Cheers,
Stan H.
=================